OpenStudy (anonymous):

Find the length of the diameter of a sphere with a volume of 321.57 miles3

OpenStudy (tkhunny):

The you DO have an idea!! Have you considered solving for "R"?

OpenStudy (anonymous):

Well that's the problem, how do I rework the problem so that I'm solving for R?

OpenStudy (tkhunny):

\(V = \dfrac{4}{3}\pi r^{3}\) \(3V = 4\pi r^{3}\) \(\dfrac{3}{4}V = \pi r^{3}\) \(\dfrac{3}{4\pi}V = r^{3}\) You're not going to make me do it all, are you?

OpenStudy (anonymous):

did i ask you to do it all? i only wanted an adjusted formula, no need to be rude.

OpenStudy (tkhunny):

Please read the Code of Conduct. It is not appropriate for me just to do the work for you. The point of this exercise is for YOU to demonstrate that YOU can manipulate the formula. I left one piece for you. Can you do it?

OpenStudy (anonymous):

Never did I ask you to do all of the work for me. There's no need for you to be condescending. In fact I'll look for help elsewhere, thanks.

OpenStudy (tkhunny):

There is neither rude nor condescending in this thread. I dare you to point out a single word that is necessarily rude or condescending. If you really want someone else to do the work for you, and no one will, I can see how you could take it that way. So, you're saying you cannot do the last part?

OpenStudy (anonymous):

Where did I say that I wanted someone to do all of the work for me? In fact, I've pointed out multiple times that I'm not. "You're not going to make me do it all, are you?" was condescending by suggesting that I was forcing you in to doing all of my work. There you go again accusing me of trying to get someone to do my work by saying "if you really want someone to do the work for you, I can see how you could take it that way". No, I'm not upset because you "won't do my work for me", I'm upset because you immediately implied that you were being forced in to doing all the work for me.

OpenStudy (tkhunny):

So, show your work! That's all I'm asking. Can you or can you not solve for \(r\)? Prove it! I left you only one step.

OpenStudy (anonymous):

First of all I don't have to prove myself to you, I really didn't appreciate your tone throughout this conversation. Second of all I didn't understand what you wrote. Took me 20 minutes but I figured it out by going to khan academy and watching like 5+ tutorials. Then I put this together and figured it out myself. V = 4/3 x Pi x r^3 r^3 = 3/4 *V/pi r^3 = 3/4 * 321.57/3.14 r^3 = 76.8 r = 4.25 (4.25 x 2) d = 8.50

OpenStudy (tkhunny):

Please read the Code of Conduct. To participate, you must show your work. If you don't like the characterization "prove it", fine. Show your work, then -- as you just did. That's the spirit. I could comment on the number of decimal places you have retained. Perhaps you are instructed to use 3.14 for \(\pi\). Generally though, good work! The more exact answer 8.5001313 is only just barely different.

OpenStudy (anonymous):

Yes I was instructed on using 3.14.

OpenStudy (tkhunny):

Perfect. Just for the record, I do get along with just about everyone if one can put up with me long enough to see that I am just honest and direct. Unfortunately, this is often misconstrued as rude or confrontational.